# Characterizing the Distinguishability of Product Distributions through Multicalibration

### v3

@@ 3.1 Proof of Part (a)
Let =$\mathcal{H}=\{h:\mathcal{X}\to[0,1]\}$ be a class of functions that is closed under negation and contains the all-zero ($h(x)=0$ for all $x$) and all-one ($h(x)=1$ for all $x$) functions. Let $\mathcal{D}$ be a distribution over $\mathcal{X}$ and consider $\varepsilon>0$. Let $\mathcal{H}^{\prime}$ be any family of functions such that $\mathcal{H}^{\prime}_{c\log|\mathcal{X}|,c}\subseteq\mathcal{H}$ for a universal constant $c$.
Suppose that $g:\mathcal{X}\to[0,1]$ is $(\mathcal{H},\varepsilon)$-indistinguishable from the constant function $v:=\mathbb{E}_{x\sim\mathcal{D}}[g(x)]$. Then the distribution $\mathcal{D}_{|g(x)=1}$ is $\left(\mathcal{H}^{\prime},\frac{\varepsilon}{v(1-v)}\right)$-indistinguishable from $\mathcal{D}_{|g(x)=0}$ for $\mathcal{H}^{\prime}$.
@@
We work out the details for showing $\mathcal{D}_{P_{i}|g(x)=1}=X_{1}|_{P_{i}}$ below, and the proof of $\mathcal{D}_{P_{i}|g(x)=0}=X_{0}|_{P_{i}}$ follows similarly. We see that
@@
$=\frac{\mathbb{P}[X_{1}\in P_{i}]}{\mathbb{P}[X_{0}\in P_{i}]+\mathbb{P}[X_{1}\in P_{i}]}.$
@@
$=\frac{1}{\mathbb{P}[X_{0}\in P_{i}]+\mathbb{P}[X_{1}\in P_{i}]}\cdot\left(\mathbb{P}[X_{0}|_{P_{i}}=x]\cdot\mathbb{P}[X_{0}\in P_{i}]+\mathbb{P}[X_{1}|_{P_{i}}=x]\cdot\mathbb{P}[X_{1}\in P_{i}]\right).$
@@
$=\sum_{x\in P_{i}}\frac{\mathbb{P}_{x\sim\mathcal{D}}[x]}{\mathbb{P}_{\mathcal{D}}[P_{i}]}\cdot\frac{\mathbb{P}[X_{1}=x]}{\mathbb{P}[X_{0}=x]+\mathbb{P}[X_{1}=x]}$
$=\frac{1}{2}\frac{1}{\mathbb{P}_{\mathcal{D}}[P_{i}]}\sum_{x\in P_{i}}\mathbb{P}[X_{1}=x]=\frac{1}{2}\frac{1}{\frac{1}{2}\mathbb{P}[X_{0}\in P_{i}]+\frac{1}{2}\mathbb{P}[X_{1}\in P_{i}]}\mathbb{P}[X_{1}\in P_{i}]$
$=\frac{\mathbb{P}[X_{1}\in P_{i}]}{\mathbb{P}[X_{0}\in P_{i}]+\mathbb{P}[X_{1}\in P_{i}]}.$
@@
$=\mathbb{P}[X_{0}\in P_{i}]\cdot\left(\alpha_{0}(P_{i})\cdot\mathbb{P}[X_{0}|_{P_{i}}=x]+\alpha_{1}(P_{i})\cdot\mathbb{P}[X_{1}|_{P_{i}}=x]\right).$
@@
$=\left|\sum_{x\in P_{i}}f(x)\cdot\left((\alpha_{0}(P_{i})-1)\mathbb{P}_{X_{0}|_{P_{i}}}[x]+\alpha_{1}(P_{i})\mathbb{P}_{X_{1}|_{P_{i}}}[x]\right)\right|$
$=\alpha_{1}(P_{i})\left|\sum_{x\in P_{i}}f(x)\cdot\left(\mathbb{P}_{X_{0}|_{P_{i}}}[x]-\mathbb{P}_{X_{1}|_{P_{i}}}[x]\right)\right|\leq\frac{\varepsilon}{1-\alpha_{1}(P_{i})}.$
@@
$+\sum_{P_{i}\in\mathcal{P}\setminus\mathcal{P}(\gamma)}\left|\sum_{x\in P_{i}}f(x)\cdot\left(\mathbb{P}[X_{0}=x]-\mathbb{P}[\widetilde{X}_{0}=x]\right)\right|.$ (6)
@@
$=\sum_{P_{i}\in\mathcal{P}(\gamma)}\mathbb{P}[X_{0}\in P_{i}]\left|\sum_{x\in P_{i}}f(x)\cdot\left(\mathbb{P}[X_{0}|_{P_{i}}=x]-\mathbb{P}[\widetilde{X}_{0}|_{P_{i}}=x]\right)\right|.$
@@
$=\sum_{P_{i}\in\mathcal{P}\setminus\mathcal{P}(\gamma)}2\mathbb{P}[X_{0}\in P_{i}]<4\gamma|\mathcal{P}|,$
@@ 3.2 Proof of Parts (b, c, d)
$p$ is computable by circuits in $\mathcal{F}_{O(1/\varepsilon^{2}),O(1/(\varepsilon^{6})\cdot\log(|\mathcal{X}|)\log(|\mathcal{X}|/\varepsilon))}$.
@@
$\mathcal{F}^{\prime}_{O(1/\varepsilon^{2}),O(1/(\varepsilon^{6})\cdot\log(|\mathcal{X}|/\varepsilon))}.$
@@
    $X_{0}^{\otimes k}$ and $X_{1}^{\otimes k}$ are $(d_{TV}(\widetilde{X}_{0}^{\otimes k},\widetilde{X}_{1}^{\otimes k})-2k\cdot\varepsilon)$-distinguishable by functions in $\mathcal{F}_{K_{2},K_{1}}$, where $K_{1}=O(\frac{k}{\varepsilon^{6}}\cdot\log(|\mathcal{X}|)\cdot\log(|\mathcal{X}|/\varepsilon))+O(k\cdot\log(|\mathcal{X}|)\cdot\mathrm{polylog}(k/\varepsilon))$ and $K_{2}=O(k/\varepsilon^{2})$.
@@ 4.1 Indistinguishability of $X_{b}^{\otimes k}$ and $\widetilde{X}_{b}^{\otimes k}$
Consider $i=0$; the case of $i=1$ follows similarly. Let $X_{0}^{\otimes(k-j)}\widetilde{X}_{0}^{\otimes j}$ denote a product distribution such that the first $k-j$ coordinates are distributed according to random variable $X_{0}$ and the remaining $j$ coordinates are distributed according to $\widetilde{X}_{0}$. Note that:
@@
$=\sum_{j=0}^{k-1}\mathbb{P}[X_{0}^{\otimes(k-j-1)}=\bar{x}_{1}\bar{x}_{2}\dots\bar{x}_{k-j-1}]\cdot\mathbb{P}[\widetilde{X}_{0}^{\otimes j}=\bar{x}_{k-j+1}\bar{x}_{k-j+2}\dots\bar{x}_{k-j-1}]\cdot\left(\mathbb{P}[X_{0}=\bar{x}_{k-j}]-\mathbb{P}[\widetilde{X}_{0}=\bar{x}_{k-j}]\right).$
@@
$-\mathbb{E}_{a\sim\widetilde{X}_{0}}\big{[}f((\bar{x}_{1},\bar{x}_{2},\dots,\bar{x}_{k-j-1},a,\bar{x}_{k-j+1},\bar{x}_{k-j+2},\dots,\bar{x}_{k}))\big{]}\bigg{|}\bigg{]}$
@@
$=\sum_{j=0}^{k-1}\mathbb{E}_{\bar{x}_{1}\bar{x}_{2}\dots\bar{x}_{k-j-1}\bar{x}_{k-j+1}\bar{x}_{k-j+2}\dots\bar{x}_{k}\sim X_{0}^{\otimes(k-j-1)}\widetilde{X}_{0}^{\otimes j}}\bigg{[}\bigg{|}\mathbb{E}_{a\sim X_{0}}\big{[}h_{\bar{x}_{1}\bar{x}_{2}\dots\bar{x}_{k-j-1}\bar{x}_{k-j+1}\bar{x}_{k-j+2}\dots\bar{x}_{k}}(a)\big{]}$
$-\mathbb{E}_{a\sim\widetilde{X}_{0}}\big{[}h_{\bar{x}_{1}\bar{x}_{2}\dots\bar{x}_{k-j-1}\bar{x}_{k-j+1}\bar{x}_{k-j+2}\dots\bar{x}_{k}}(a)\big{]}\bigg{|}\bigg{]}.$
@@
$=O(k\varepsilon)+\sqrt{2}\sqrt{1-(1-d_{H}^{2}(\widetilde{X}_{0},\widetilde{X}_{1}))^{k}}.$
@@ 4.3 Small Circuits that Distinguish Between $X_{0}^{\otimes k},X_{1}^{\otimes k}$: Proof of Item (3)
with $K_{1}=k/\varepsilon^{6}\cdot\log(|\mathcal{X}|)\cdot\log(|\mathcal{X}|/\varepsilon)+O(k\cdot\mathrm{polylog}(1/(\varepsilon\delta)))$, and $K_{2}=O(k/\varepsilon^{2})$ which distinguishes $X^{\otimes k}_{0},X^{\otimes k}_{1}$ with advantage $d_{\mathrm{TV}}(\widetilde{X}_{0}^{\otimes k},\widetilde{X}_{1}^{\otimes k})-4\delta k$.
@@
For random variables $X_{0},X_{1}$ and family of functions $\mathcal{F}$, consider the random variables $\widetilde{X}_{0},\widetilde{X}_{1}$, and partition function $p$ as defined in Theorem 3.1. Recall that by this theorem, the partitions $\mathcal{P}\in\mathcal{F}_{q,t}$, with $t=O(1/(\varepsilon^{6})\cdot\log(|\mathcal{X}|)/\varepsilon),q=O(1/\varepsilon^{2})$, in the sense that there is a circuit of the above stated complexity which can compute the index of the partition that each element is in.
@@
For a sequence of $k$ samples $x_{1},\dots x_{k}$, the optimal distinguisher is simply the distinguisher which returns $1$ (corresponding to deciding that the samples are from $\widetilde{X}_{1,\delta}^{\otimes k}$) if the above expression is $\geq 1$, and 00 otherwise (corresponding to deciding that the samples are from $\widetilde{X}_{0,\delta}^{\otimes k}$). This is because in general for two distributions $p,q$, we have that
@@
Now, we will show that we can actually compute such a circuit $C$ which is not too complex. Indeed, the key observation is that because we “rounded” our distributions initially, now when we compute the product $\prod_{j=1}^{k}\frac{\alpha_{1,\delta}(P_{i_{j}})}{\alpha_{0,\delta}(P_{i_{j}})}$, every pair of $\alpha_{1,\delta}(P_{i_{j}}),\alpha_{0,\delta}(P_{i_{j}})$ share a denominator. Thus, computing this product is equivalent to multiplying together $k$ integers in $[r/\delta]$. So, we consider the following circuit: first, $C$ has $k$ copies of the circuit which, upon receiving element $x$, computes the index $i$ for which $x\in P_{i}$. Recall that this circuit is $\in\mathcal{F}_{t,q}$. Next, once we have computed the index $i$ of the partition that an element $x$ is in, we hard-wire two-lookup tables: one computes the numerator of $\alpha_{0,\delta}(P_{i})$, and the other computes the numerator of $\alpha_{1,\delta}(P_{i})$. Observe that because there are only $r$ parts in the partition, the complexity of these look-up tables is not too large. By [Lup70], each look-up table requires only circuits of size $O(r)$. Once we have computed the numerators of these values, all that remains is to multiply together these integers, and return the index $b\in\{0,1\}$, corresponding to whichever product is larger. Computing this product can be done with circuits of size $O(k\cdot\mathrm{polylog}(r/\delta))$.
@@
as we use $k$ copies of a circuit in $\mathcal{F}_{q,t}$, along with $O(k)$ look-up tables, and $O(k\cdot\mathrm{polylog}(r/\delta))$ extra wires to compute the product. By plugging in $q=O(1/\varepsilon^{2}),r=O(1/\varepsilon),t=O(1/(\varepsilon^{6}\gamma)\cdot\log(|\mathcal{X}|)\cdot\log(|\mathcal{X}|)/\varepsilon)$, we obtain the stated complexity.
@@
$=d_{\mathrm{TV}}(\widetilde{X}_{0,\delta}^{\otimes k},\widetilde{X}_{1,\delta}^{\otimes k})\geq d_{\mathrm{TV}}(\widetilde{X}_{0}^{\otimes k},\widetilde{X}_{1}^{\otimes k})-2\delta k.$
@@
$-\left|\sum_{(x_{1},\dots x_{k})\in\mathcal{X}^{k}}C((x_{1},\dots x_{k}))\cdot\left[\widetilde{X}_{0}^{\otimes k}(x_{1},\dots x_{k})-\widetilde{X}_{0,\delta}^{\otimes k}(x_{1},\dots x_{k})\right]\right|$
$-\left|\sum_{(x_{1},\dots x_{k})\in\mathcal{X}^{k}}C((x_{1},\dots x_{k}))\cdot\left[\widetilde{X}_{1}^{\otimes k}(x_{1},\dots x_{k})-\widetilde{X}_{1,\delta}^{\otimes k}(x_{1},\dots x_{k})\right]\right|$
@@
$-d_{\mathrm{TV}}(\widetilde{X}_{0}^{\otimes k},\widetilde{X}_{0,\delta}^{\otimes k})-d_{\mathrm{TV}}(\widetilde{X}_{1}^{\otimes k},\widetilde{X}_{1,\delta}^{\otimes k})$
@@
$=\left|\sum_{(x_{1},\dots x_{k})\in\mathcal{X}^{k}}C((x_{1},\dots x_{k}))\cdot\left[X_{0}^{\otimes k}(x_{1},\dots x_{k})-X_{1}^{\otimes k}(x_{1},\dots x_{k})\right]\right|.$

### v5

@@ 3.1 Proof of Part (a)
Let $\mathcal{H}=\{h:\mathcal{X}\to[0,1]\}$ be a class of functions that is closed under negation and contains the all-zero ($h(x)=0$ for all $x$) and all-one ($h(x)=1$ for all $x$) functions. Let $\mathcal{D}$ be a distribution over $\mathcal{X}$ and consider $\varepsilon>0$. Let $\mathcal{H}^{\prime}$ be any family of functions such that $\mathcal{H}^{\prime}_{c\log|\mathcal{X}|,c}\subseteq\mathcal{H}$ for a universal constant $c$.
Suppose that $g:\mathcal{X}\to[0,1]$ is identified with a randomized Boolean-valued function and is $(\mathcal{H},\varepsilon)$-indistinguishable from the constant function $v:=\mathbb{E}_{x\sim\mathcal{D}}[g(x)]$. Then the distribution $\mathcal{D}_{|g(x)=1}$ is $\left(\mathcal{H}^{\prime},\frac{\varepsilon}{v(1-v)}\right)$-indistinguishable from $\mathcal{D}_{|g(x)=0}$ for $\mathcal{H}^{\prime}$.
@@
We work out the details for showing $\mathcal{D}_{P_{i}|g(x)=1}=X_{1}|_{P_{i}}$ below, and the proof of $\mathcal{D}_{P_{i}|g(x)=0}=X_{0}|_{P_{i}}$ follows similarly. In what follows, let $\mathbb{P}_{\text{rand}(g)}$ denote that the probability of an event is taken over the randomness of the randomized Boolean-valued function $g$.
We see that
@@
$\frac{\mathbb{P}[X_{1}\in P_{i}]}{\mathbb{P}[X_{0}\in P_{i}]+\mathbb{P}[X_{1}\in P_{i}]}.$
@@
$\frac{1}{\mathbb{P}[X_{0}\in P_{i}]+\mathbb{P}[X_{1}\in P_{i}]}\cdot\left(\mathbb{P}[X_{0}|_{P_{i}}=x]\cdot\mathbb{P}[X_{0}\in P_{i}]+\mathbb{P}[X_{1}|_{P_{i}}=x]\cdot\mathbb{P}[X_{1}\in P_{i}]\right).$
@@
$\sum_{x\in P_{i}}\frac{\mathbb{P}_{x\sim\mathcal{D}}[x]}{\mathbb{P}_{\mathcal{D}}[P_{i}]}\cdot\frac{\mathbb{P}[X_{1}=x]}{\mathbb{P}[X_{0}=x]+\mathbb{P}[X_{1}=x]}$
$\frac{1}{2}\frac{1}{\mathbb{P}_{\mathcal{D}}[P_{i}]}\sum_{x\in P_{i}}\mathbb{P}[X_{1}=x]=\frac{1}{2}\frac{1}{\frac{1}{2}\mathbb{P}[X_{0}\in P_{i}]+\frac{1}{2}\mathbb{P}[X_{1}\in P_{i}]}\mathbb{P}[X_{1}\in P_{i}]$
$\frac{\mathbb{P}[X_{1}\in P_{i}]}{\mathbb{P}[X_{0}\in P_{i}]+\mathbb{P}[X_{1}\in P_{i}]}.$
@@
$\mathbb{P}[X_{0}\in P_{i}]\cdot\left(\alpha_{0}(P_{i})\cdot\mathbb{P}[X_{0}|_{P_{i}}=x]+\alpha_{1}(P_{i})\cdot\mathbb{P}[X_{1}|_{P_{i}}=x]\right).$
@@
$\left|\sum_{x\in P_{i}}f(x)\cdot\left((\alpha_{0}(P_{i})-1)\mathbb{P}_{X_{0}|_{P_{i}}}[x]+\alpha_{1}(P_{i})\mathbb{P}_{X_{1}|_{P_{i}}}[x]\right)\right|$
$\alpha_{1}(P_{i})\left|\sum_{x\in P_{i}}f(x)\cdot\left(\mathbb{P}_{X_{0}|_{P_{i}}}[x]-\mathbb{P}_{X_{1}|_{P_{i}}}[x]\right)\right|\leq\frac{\varepsilon}{1-\alpha_{1}(P_{i})}.$
@@
$\sum_{P_{i}\in\mathcal{P}\setminus\mathcal{P}(\gamma)}\left|\sum_{x\in P_{i}}f(x)\cdot\left(\mathbb{P}[X_{0}=x]-\mathbb{P}[\widetilde{X}_{0}=x]\right)\right|.$ (6)
@@
$\sum_{P_{i}\in\mathcal{P}(\gamma)}\mathbb{P}[X_{0}\in P_{i}]\left|\sum_{x\in P_{i}}f(x)\cdot\left(\mathbb{P}[X_{0}|_{P_{i}}=x]-\mathbb{P}[\widetilde{X}_{0}|_{P_{i}}=x]\right)\right|.$
@@
$\sum_{P_{i}\in\mathcal{P}\setminus\mathcal{P}(\gamma)}2\mathbb{P}[X_{0}\in P_{i}]<4\gamma|\mathcal{P}|,$
@@ 3.2 Proof of Parts (b, c, d)
$p$ is computable by circuits in $\mathcal{F}_{O(1/\varepsilon^{6}),O(1/(\varepsilon^{12})\cdot\log(|\mathcal{X}|)\log(|\mathcal{X}|/\varepsilon))}$.
@@
$\mathcal{F}^{\prime}_{O(1/\varepsilon^{6}),O(1/(\varepsilon^{12})\cdot\log(|\mathcal{X}|/\varepsilon))}.$
@@
    $X_{0}^{\otimes k}$ and $X_{1}^{\otimes k}$ are $(d_{TV}(\widetilde{X}_{0}^{\otimes k},\widetilde{X}_{1}^{\otimes k})-2k\cdot\varepsilon)$-distinguishable by functions in $\mathcal{F}_{K_{2},K_{1}}$, where $K_{1}=O(\frac{k}{\varepsilon^{12}}\cdot\log(|\mathcal{X}|)\cdot\log(|\mathcal{X}|/\varepsilon))+O(k\cdot\log(|\mathcal{X}|)\cdot\mathrm{polylog}(k/\varepsilon))$ and $K_{2}=O(k/\varepsilon^{6})$.
@@ 4.1 Indistinguishability of $X_{b}^{\otimes k}$ and $\widetilde{X}_{b}^{\otimes k}$
Consider $i=0$; the case of $i=1$ follows similarly. Let $X_{0}^{\otimes(k-j)}\widetilde{X}_{0}^{\otimes j}$ denote a product distribution such that the first kjk-j coordinates are distributed according to random variable $X_{0}$ and the remaining $j$ coordinates are distributed according to $\widetilde{X}_{0}$. Note that:
@@
$\sum_{j=0}^{k-1}\mathbb{P}[X_{0}^{\otimes(k-j-1)}=\bar{x}_{1}\bar{x}_{2}\dots\bar{x}_{k-j-1}]\cdot\mathbb{P}[\widetilde{X}_{0}^{\otimes j}=\bar{x}_{k-j+1}\bar{x}_{k-j+2}\dots\bar{x}_{k-j-1}]\cdot\left(\mathbb{P}[X_{0}=\bar{x}_{k-j}]-\mathbb{P}[\widetilde{X}_{0}=\bar{x}_{k-j}]\right).$
@@
$\mathbb{E}_{a\sim\widetilde{X}_{0}}\big[f((\bar{x}_{1},\bar{x}_{2},\dots,\bar{x}_{k-j-1},a,\bar{x}_{k-j+1},\bar{x}_{k-j+2},\dots,\bar{x}_{k}))\big]\bigg|\bigg]$
@@
$\sum_{j=0}^{k-1}\mathbb{E}_{\bar{x}_{1}\bar{x}_{2}\dots\bar{x}_{k-j-1}\bar{x}_{k-j+1}\bar{x}_{k-j+2}\dots\bar{x}_{k}\sim X_{0}^{\otimes(k-j-1)}\widetilde{X}_{0}^{\otimes j}}\bigg[\bigg|\mathbb{E}_{a\sim X_{0}}\big[h_{\bar{x}_{1}\bar{x}_{2}\dots\bar{x}_{k-j-1}\bar{x}_{k-j+1}\bar{x}_{k-j+2}\dots\bar{x}_{k}}(a)\big]$
$\mathbb{E}_{a\sim\widetilde{X}_{0}}\big[h_{\bar{x}_{1}\bar{x}_{2}\dots\bar{x}_{k-j-1}\bar{x}_{k-j+1}\bar{x}_{k-j+2}\dots\bar{x}_{k}}(a)\big]\bigg|\bigg].$
@@
$O(k\varepsilon)+\sqrt{2}\sqrt{1-(1-d_{H}^{2}(\widetilde{X}_{0},\widetilde{X}_{1}))^{k}}.$
@@ 4.3 Small Circuits that Distinguish Between $X_{0}^{\otimes k},X_{1}^{\otimes k}$: Proof of Item (3)
with $K_{1}=k/\varepsilon^{12}\cdot\log(|\mathcal{X}|)\cdot\log(|\mathcal{X}|/\varepsilon)+O(k\cdot\mathrm{polylog}(1/(\varepsilon\delta)))$, and $K_{2}=O(k/\varepsilon^{6})$ which distinguishes $X^{\otimes k}_{0},X^{\otimes k}_{1}$ with advantage $d_{\mathrm{TV}}(\widetilde{X}_{0}^{\otimes k},\widetilde{X}_{1}^{\otimes k})-4\delta k$.
@@
For random variables $X_{0},X_{1}$ and family of functions $\mathcal{F}$, consider the random variables $\widetilde{X}_{0},\widetilde{X}_{1}$, and partition function $p$ as defined in Theorem 3.1. Recall that by this theorem, the partitions $\mathcal{P}\in\mathcal{F}_{q,t}$, with $t=O(1/(\varepsilon^{6})\cdot\log(|\mathcal{X}|)/\varepsilon),q=O(1/\varepsilon^{6})$, in the sense that there is a circuit of the above stated complexity which can compute the index of the partition that each element is in.
@@
For a sequence of $k$ samples $x_{1},\dots x_{k}$, the optimal distinguisher is simply the distinguisher which returns $1$ (corresponding to deciding that the samples are from $\widetilde{X}_{1,\delta}^{\otimes k}$) if the above expression is $\geq 1$, and 0 otherwise (corresponding to deciding that the samples are from $\widetilde{X}_{0,\delta}^{\otimes k}$). This is because in general for two distributions $p,q$, we have that
@@
Now, we will show that we can actually compute such a circuit $C$ which is not too complex. Indeed, the key observation is that because we “rounded” our distributions initially, now when we compute the product $\prod_{j=1}^{k}\frac{\alpha_{1,\delta}(P_{i_{j}})}{\alpha_{0,\delta}(P_{i_{j}})}$, every pair of $\alpha_{1,\delta}(P_{i_{j}}),\alpha_{0,\delta}(P_{i_{j}})$ share a denominator. Thus, computing this product is equivalent to multiplying together $k$ integers in $r/\delta]$. So, we consider the following circuit: first, $C$ has $k$ copies of the circuit which, upon receiving element $x$, computes the index $i$ for which $x\in P_{i}$. Recall that this circuit is $\in\mathcal{F}_{t,q}$. Next, once we have computed the index $i$ of the partition that an element $x$ is in, we hard-wire two-lookup tables: one computes the numerator of $\alpha_{0,\delta}(P_{i})$, and the other computes the numerator of $\alpha_{1,\delta}(P_{i})$. Observe that because there are only $r$ parts in the partition, the complexity of these look-up tables is not too large. By [Lup70], each look-up table requires only circuits of size $O(r)$. Once we have computed the numerators of these values, all that remains is to multiply together these integers, and return the index $b\in\{0,1\}$, corresponding to whichever product is larger. Computing this product can be done with circuits of size $O(k\cdot\mathrm{polylog}(r/\delta))$.
@@
as we use $k$ copies of a circuit in $\mathcal{F}_{q,t}$, along with $O(k)$ look-up tables, and $O(k\cdot\mathrm{polylog}(r/\delta))$ extra wires to compute the product. By plugging in $q=O(1/\varepsilon^{6}),r=O(1/\varepsilon),t=O(1/(\varepsilon^{6}\gamma)\cdot\log(|\mathcal{X}|)\cdot\log(|\mathcal{X}|)/\varepsilon)$, we obtain the stated complexity.
@@
$d_{\mathrm{TV}}(\widetilde{X}_{0,\delta}^{\otimes k},\widetilde{X}_{1,\delta}^{\otimes k})\geq d_{\mathrm{TV}}(\widetilde{X}_{0}^{\otimes k},\widetilde{X}_{1}^{\otimes k})-2\delta k.$
@@
$\left|\sum_{(x_{1},\dots x_{k})\in\mathcal{X}^{k}}C((x_{1},\dots x_{k}))\cdot\left[\widetilde{X}_{0}^{\otimes k}(x_{1},\dots x_{k})-\widetilde{X}_{0,\delta}^{\otimes k}(x_{1},\dots x_{k})\right]\right|$
$\left|\sum_{(x_{1},\dots x_{k})\in\mathcal{X}^{k}}C((x_{1},\dots x_{k}))\cdot\left[\widetilde{X}_{1}^{\otimes k}(x_{1},\dots x_{k})-\widetilde{X}_{1,\delta}^{\otimes k}(x_{1},\dots x_{k})\right]\right|$
@@
$d_{\mathrm{TV}}(\widetilde{X}_{0}^{\otimes k},\widetilde{X}_{0,\delta}^{\otimes k})-d_{\mathrm{TV}}(\widetilde{X}_{1}^{\otimes k},\widetilde{X}_{1,\delta}^{\otimes k})$
@@
$\left|\sum_{(x_{1},\dots x_{k})\in\mathcal{X}^{k}}C((x_{1},\dots x_{k}))\cdot\left[X_{0}^{\otimes k}(x_{1},\dots x_{k})-X_{1}^{\otimes k}(x_{1},\dots x_{k})\right]\right|.$